OpenStudy (anonymous):

A square loop of wire, 15 cm on a side, carries a current of 0.50 A. It stands vertically, such that its normal is parallel to a uniform horizontal magnetic field that exists in a region in which half of the loop is immersed, with the other half below the region. The mass per unit length of the wire is 5.6 g/m. What is the strength of the magnetic field if the loop simply hangs there?

OpenStudy (perl):

this is physics , i need you equations

OpenStudy (anonymous):

yes it's physics, but in physics part there is no one help in anything :S

OpenStudy (perl):

sorry dont know the flux equation

OpenStudy (perl):

is it an integral??

OpenStudy (anonymous):

it's ok :)

OpenStudy (perl):

i guess you might have luck here, do you know a website that has this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

mcgraw-hill.com

OpenStudy (perl):

Once the switch is closed, the line voltage is applied across the bulb. As the voltage is applied across the cold filament when it is first turned on, the resistance of the filament is low, the current is high, and a relatively large amount of power is delivered to the bulb. As the filament warms, its resistance rises and the current decreases. As a result, the power delivered to the bulb decreases. The large current spike at the beginning of the bulb’s operation is the reason that lightbulbs often fail just after they are turned on.

OpenStudy (perl):

what page bro?

OpenStudy (perl):

that was interesting fact i found when i searched your question

OpenStudy (anonymous):

this question from assignment

OpenStudy (anonymous):

so i didn't know what page but it relate to ebook on this site

OpenStudy (perl):

they have a whole physics book?? online , wow

OpenStudy (anonymous):

yes, but I am confused now because the deadline of this assignment after 4 hours

OpenStudy (anonymous):

for*

OpenStudy (anonymous):

BIL= mg or B*0.50*0.15= 0.0056*0.15*4*9.81 or B = (0.0056*4*9.81)/0.50 = 0.439 or 0.44 T